OpenStudy (anonymous):

a box contains 95 pink rubber band and 90 brown rubber bands. you select a rubber band at random from the box. find each probability. write the probability as a fraction in simplest form. A. find the theoretical probability of selecting a pink rubber band b. find the theoretical probability of selecting a brown rubber band

OpenStudy (rulnick):

P(pink) = 95 / (95+90) = 95/185 = ... (reduce the fraction) P(brown) = 90 / 185 = ... (reduce) ok?

OpenStudy (rulnick):

95/185 reduced is 19/37 (divide numerator and denominator by 5)

OpenStudy (rulnick):

Likewise, 90/185 can be reduced to ?/37.

OpenStudy (anonymous):

what bout 95/185

OpenStudy (rulnick):

You mean 90/185? It's something over 37. To get the something, divide 90 by 5.

OpenStudy (anonymous):

do i do the same for 95/185

OpenStudy (rulnick):

Yes, but I did that one for you (above).

OpenStudy (anonymous):

but will i divide by 5 for the other one